My son's teacher marked these wrong in his Math test. My understanding is she expected them to be written as subtraction equations even though it was not specifically mentioned in the test. So please help me understand.
1)
Mr. Pickles owes $350 on his laptop. He has already paid $930 on it. What equation models determining the original cost of the laptop.
This is my son's equation which was marked as incorrect :

350 + 930 = c ; c = 1280
2) Katy has to sell some Girl scout cookies to her friends.If she sells 9 dozen she will have 24 dozen left. How many dozens does she have to begin with?
This is my son's equation which was marked as incorrect :

9 + 24 = d ; d = 33 dozen
What instructions in both questions points to them being subtraction equations ?
Would greatly appreciate your help.

Found 2 solutions by josgarithmetic, math_helper:
Answer by josgarithmetic(39617)

---------
---------
1)
Mr. Pickles owes $350 on his laptop. He has already paid $930 on it. What equation models determining the original cost of the laptop.
This is my son's equation which was marked as incorrect :
350 + 930 = c ; c = 1280
---------
---------

Let c be the original cost (or price) of the laptop.
If Mr. P still OWES $350, and already paid $930, then an equation using subtraction can be: c-930=350.
Also true can be this one, showing what what already paid: c-350=930.

Answer by math_helper(2461)

Solving b/c the student asked me.
1) c = original cost
c - 930 = 350 (from "he has already paid $930...")


2) d = original number of dozens
d - 9 = 24 (from "If she sells 9 dozen...")

-----
In solving algebra word problems (and many real world problems) there are two distinct steps:
(1) Model the problem. This means writing an equation that precisely represents the real world problem.

(2) Perform the algebraic manipulations to solve for the unknown(s) (down the road there will be more than one unknown for some problems). Sometimes this step is called "turning the crank."

This teacher is checking if the student can model the problem before turning the crank.